RC Meeting and Run Plan, Friday, February 14th 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

Friday 14th Feb. 2020.

General updates

  • During Thursday shifts:
    • Finished data taking with reference cell and carbon optics at high and low momentum settings except for carbon optics at low momentum.
    • Took data with 3He transverse about 1 hr.
  • Today:
    • Wien flip for Hall A and pass change for Hall B are postponed to Monday. SLM replacement in South Linac.
    • Power supply swap (Bill) and reference cell pressure calibration (Brad) done during the down time.

Pending problem

  • Communication with MCC.

Fixed problem

Recent polarization measurements

  • Transverse: 49.62% [1]
  • Longitudinal: 49.66% [2]

Current Run Plan

    • Data taking with carbon optics at lower momentum setting (done)
      • Momentum settings: SHMS=-2.6 GeV, HMS=-2.9 GeV
      • 0.5h beam-on time (50mC charge goal) on "Carbon Optics" (not Carbon Foil, not Carbon Hole), 40uA with raster 0.5mm diameter
    • Take production data when the target is ready

  • If beam goes away for more than 2h, or if during the downtime, MCC does operations that may change the beam tuning (except work on transport magnets), then do beam recovery procedure (do YAG viewer part only if your are suspicious of the harp scan results)
  • Make sure all the beam specs are followed (e.g. no raster size below 4mm on 3He cell).
  • Always take 120Hz of EDTM
  • Fill BTA with realistic numbers. You may follow the guidance coming from "load from EPICS", if it does not overestimate the ABU (as it generally does).
!!                                                     !!
!!  ATTN: SHIFT LEADERS                                !!
!!     Please follow the shift summary template on     !!
!!     the laminated SL instructions.                  !!
!!                                                     !!
!!  ATTN: TARGET OPERATORS                             !!
!!  All NMR measurement HCLOG entries should include:  !!
!!     1) A title that clearly lists polarization      !!
!!     2) Screenshots of both NMR_FieldSweep and       !!
!!        the polarization fit                         !!
!!                                                     !!
!!     For an example see the following entry:         !!
!!     https://logbooks.jlab.org/entry/3768696       !!
!!                                                     !!


  • General idea: if target polarization is about equal for transverse vs. longitudinal, alternate 4 hours of transverse and 20 hours of longitudinal data taking. Do NMR before and after transverse, and every 5 hours during longitudinal.
  • Production at 30 degrees. We may want to start from HMS/SHMS=-2.9/-2.6 momentum setting, complete this and then go back to HMS/SHMS=-3.5/-3.4 momentum.

Hall Access

  • The next possible hall access is on Friday, Feb 14th, 2020 when the Wien flip is scheduled by the accelerator.
    • Hall C safety walkthrough (5 people)
    • Installation of the spare rotation stage. Are the mounting materials ready?
    • Power supply swap - replace Agilient with new Kepco.

When there is a down:

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
    • If SHMS dipole trips again, take a screenshot of the PSU window before resetting the interlock and post it to hclog.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • The operators have been somewhat careful with our beam orbit locks. Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[3]
  • Temporary Laser instructions can be found here [4]
  • Do not change SHMS momentum until we fixed the current lead valve

Planned Accelerator Configuration Changes

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links